OpenStudy (anonymous):

Evaluate 12 ÷ (3 + 9)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

do 9=3+12 12/12=1

OpenStudy (anonymous):

answer is 1

OpenStudy (anonymous):

Use Bodmas Bo- brackets off that is solve what is in the brackets d- division m- multiplication a- addition s- subtration

OpenStudy (maeganr):

12 divided by ( 3 + 9 ) = 1 You do the order of PEMDAS. Parenthesis: 3+9=12 || 12 divided by 12 is 1. (:

OpenStudy (anonymous):

or use Please- parenthesis Excuse- exponents My-multiplication Dear-division Aunt-adding Sally-subtracting